OpenStudy (anonymous):

Determine the LCM of the given polynomials: Leave your answer in factored form X^2+10x+25 x^2+12x+35

OpenStudy (anonymous):

Factor both first.

OpenStudy (anonymous):

(x+5)(x+5) and (x+5)(x+7)

OpenStudy (anonymous):

Ok, you need something that is a multiple of both, so you'll need both (x+5)'s from the first one. The second has a (x+5), but you already have plenty of those, so just take the (x+7) from it.

OpenStudy (anonymous):

so (x+7)

OpenStudy (anonymous):

From the second one yes, because you already have (x+5) from the first polynomial. Remember, you are looking for a multiple of both, so you have to include all the factors of both.

OpenStudy (anonymous):

It's the same as if you were looking for the LCM of 25 and 35. 25=5*5, 35=5*7, so the LCM is 5*5*7.